### v1

@@ 3.1. Proper Convex Cones
Using Lemma 2.4 we show that ${\mathcal{K}}$-nonnegativity is a sufficient condition for ${\mathcal{K}}$-Lorentzian polynomials if ${\mathcal{K}}$ is contained in its dual cone, ${\mathcal{K}}^{\ast}$.
@@
Since $Q$ is ${\mathcal{K}}$-nonnegative, so $Q({\mathcal{K}})\subseteq{\mathcal{K}}\subseteq{\mathcal{K}}^{\ast}$. Then the rest follows from Lemma 2.4. ∎
Based on Lemma 2.4 we show that ${\mathcal{K}}$-nonnegativity is a necessary condition for ${\mathcal{K}}$-Lorentzian polynomials if ${\mathcal{K}}$ contains its dual cone, ${\mathcal{K}}^{\ast}$.
@@
By Lemma 2.4 we have the matrix $Q$ of $q$ has exactly one positive eigenvalue and $Q({\mathcal{K}})\subseteq{{\mathcal{K}}^{*}}$. Thus, $Q({\mathcal{K}})\subseteq{\mathcal{K}}^{*}\subseteq{\mathcal{K}}$. Therefore, $Q$ is ${\mathcal{K}}$-nonnegative by Definition 2.3. ∎
@@
The converse to Corollary 3.2 is not true. It’s easy to construct counterexamples where ${\mathcal{K}}^{\ast}\subseteq{\mathcal{K}}$ and $Q({\mathcal{K}})\subseteq{\mathcal{K}}$ but $Q({\mathcal{K}})\not\subseteq{\mathcal{K}}^{\ast}$.
It’s shown that any closed convex cone contains its dual cone, i.e., ${\mathcal{K}}^{\ast}\subset{\mathcal{K}}$, if and only if there exist $y,z\in{\mathcal{K}}$ such that $x=y-z$ and $\langle y,z\rangle=0$, cf. [HH69] for detail. This is used to show that every $x\in{\mathbb{R}}^{n}$ has an (unique) orthogonal decomposition on ${\mathcal{K}}$ in [SV70] which have many well-known consequences. Additionally, the sum of two ${\mathcal{K}}$-nonnegative matrices is ${\mathcal{K}}$ nonnegative. Furthermore, if $A$ and $B$ are ${\mathcal{K}}$-nonnegative and $A$ is ${\mathcal{K}}$-irreducible, then so is $A+B$, cf. [BP94, Chap-1]. However, the sum of two log-concave polynomials may not be log-concave. The sufficient condition for the sum of two ${\mathcal{K}}$-Lorentzian polynomials to also be ${\mathcal{K}}$-Lorentzian can be found in [ALOGV24, Lemma 3.3] for the nonnegative orthant, and [BD24, Theorem 4.15] for any proper convex cone ${\mathcal{K}}$.
@@
[BD24, Theorem 4.15] Let $f,g\in{\mathbb{R}}^{d}_{n}[x]$ be two ${\mathcal{K}}$-Lorentzian polynomials over the self-dual cone ${\mathcal{K}}$. Then the sum $f+g$ is ${\mathcal{K}}$-Lorentzian if there exist vectors $b,c\in{\mathcal{K}}$ such that $D_{b}f=D_{c}g\not\equiv 0$.
@@ 3.2. Perron-Frobenius Theorem for Cones
Let ${\mathcal{K}}$ be a proper convex cone. We can leverage the relationship between ${\mathcal{K}}$-Lorentzian polynomials and the generalized Perron-Frobenius theorem [KR48], [BS75] to distinguish potential ${\mathcal{K}}$-Lorentzian polynomial candidates from given log-concave polynomials. This also provides valuable insight into the structure of the cone K.
@@
(Generalization of Perron-Frobenius Theorem) [KR48], [BS75]: Let $A$ be a ${\mathcal{K}}$-irreducible matrix with spectral radius $\rho$. Then
@@
By applying [BD24, Lemma 4.12] and [BD24, Prop:4.13] we derive the following characterization of the Hessian matrices evaluated at any point over the cone ${\mathcal{K}}$ for any ${\mathcal{K}}$-Lorentzian polynomials where ${\mathcal{K}}$ is a proper convex cone.
@@
In special cases using the Definition 2.2 and [BD24, Them:5.1] we have some necessary conditions for ${\mathcal{K}}$-Lorentzian polynomials over the self-dual cone ${\mathcal{K}}$.
@@
By Theorem 3.5 we have more stronger necessary conditions for ${\mathcal{K}}$-Lorentzian polynomials over the self-dual cone ${\mathcal{K}}$.
@@
Since $f$ is a ${\mathcal{K}}$-Lorentzian polynomial, so for any $a\in{\mathcal{K}}$, the nonsingular Hessian $H_{f}(a)$ is ${\mathcal{K}}$-irreducible matrix and has exactly one positive eigenvalue. Therefore, if $H_{f}(a)$ is nonsingular, by Theorem 3.5 this positive eigenvalue necessarily corresponds to the spectral radius, i.e., the positive eigenvalue must be the peripheral eigenvalue of $H_{f}(a)$, and its corresponding eigenvector must lie in the interior of the cone ${\mathcal{K}}$. ∎
The geometry of the set of symmetric Perron–Frobenius matrices has been studied in [Tar18]. In fact, it is not convex but is star convex with the identity matrix as the center. It’s a natural question to describe the geometry of the set of symmetric Perron–Frobenius matrices which have exactly one positive eigenvalue. We leave this question for future work.
@@ 3.3. Properties of ${\mathcal{K}}$-CLC polynomials
Here we propose another definition for ${\mathcal{K}}$-CLC multivariate polynomials by restricting them on certain lines. This allows us to show the log-concavity property among the coefficients of the corresponding univariate restrictions of a ${\mathcal{K}}$-CLC polynomial. Recall that a univariate polynomial $f(t)$ is log-concave over an interval $I$ if $f(t)>0$ for all $t\in I$ and $f(t)f^{{}^{\prime\prime}}(t)\leq(f^{{}^{\prime}}(t))^{2}$ for all $t\in I$. Equivalently, a univariate polynomial $f(t)$ is log-concave over an interval $I$ if $f(t)>0$ for all $t\in I$ and $\frac{d^{2}}{dt^{2}}\log f(t)\leq 0$ for all $t\in I$. So, we can define ${\mathcal{K}}$-CLC for any real multivariate polynomial as in Definition 2.1, need not be homogeneous. If they are homogeneous, the set of ${\mathcal{K}}$-Lorentzian polynomials of degree $d$ are equivalent to the set of ${\mathcal{K}}$-CLC forms , cf. [BD24]). On the other hand, over the nonnegative orthant, a characterization for bivariate ${\mathcal{K}}$-CLC form has been established in the literature, providing a resolution to Mason’s conjecture.
@@
[Gur09], [ALOGV24] $f=\sum_{k=0}^{n}c_{k}x^{n-k}y^{k}\in{\mathbb{R}}[x,y]$ is completely log- concave if and only if the sequence of nonnegative coefficients, $\{c_{0},\dots,c_{n}\}$ is ultra log-concave, i.e., for every $1<k<n$, $\left(\frac{c_{k}}{\binom{n}{k}}\right)^{2}\geq\frac{c_{k-1}}{\binom{n}{k-1}}\frac{c_{k+1}}{\binom{n}{k+1}}$.
@@
It’s straightforward to verify that $f(t)=\sum_{i=0}^{d}a_{i}t^{i}$ with $a_{i}>0$ is CLC for $t\geq 0$ if and only if its homogenization $f(t,u)=\frac{u^{d}}{d!}a_{0}+\frac{u^{d-1}}{(d-1)!}a_{1}t+\dots+\frac{u^{0}}{0!}a_{d}t^{d}$ is CLC over ${\mathbb{R}}^{2}_{\geq 0}$. Then the rest follows from Theorem 3.9 and the identity $a_{i}^{2}(i!)^{2}\geq a_{i-1}a_{i+1}(i-1)!(i+1)!\Leftrightarrow ia_{i}^{2}\geq(i+1)a_{i-1}a_{i+1}$. ∎
@@
Due to Lemma 3.10 and symmetry, $((d-k)!a_{k})^{2}\geq(d-k-1)!(d-k+1)!a_{k-1}a_{k+1}\Leftrightarrow(d-k)a_{k}^{2}\geq(d-k+1)a_{k-1}a_{k+1}\Leftrightarrow ia_{d-i}^{2}\geq(i+1)a_{d-i-1}a_{d-i+1}$. ∎
We discuss some properties of ${\mathcal{K}}$-CLC polynomials which lead to another characterization of ${\mathcal{K}}$CLC in terms of its line restriction. It is important to note that the convexity of a function can be characterized in terms of one-dimensional convexity as follows. A function $f:{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}$ is convex over a convex set ${\mathcal{K}}$ if and only if for any $x\in{\mathcal{K}}$ and $v\in{\mathbb{R}}^{n}$, the function $g:{\mathbb{R}}\rightarrow{\mathbb{R}},g(t)=f(x+tv)$ is convex as function in $t$ for all $t$ such that $x+tv\in{\mathcal{K}}$ [BV04]. In the same spirit, we introduce a definition of ${\mathcal{K}}$-CLC polynomial after restricting it to certain lines.
Definition 3.12.
$f(x)\in{\mathbb{R}}[x]$ is (completely) log-concave over a proper convex cone ${\mathcal{K}}\subseteq{\mathbb{R}}^{n}$ if and only if for any $x\in{\mathcal{K}}$ and $v\in{\mathbb{R}}^{n}$, $g(t)=f(x+tv)$ is (completely) log-concave for all $t$ such that $x+tv\in{\mathcal{K}}$.
@@
Let $f(x)$ be a strictly ${\mathcal{K}}$-CLC over a proper convex cone ${\mathcal{K}}$. Then for any $x,v\in{\mathcal{K}}$, the coefficients of $f(x+tv)$ are positive.
@@
It follows from Definition 2.1 and Proposition 3.13. ∎
@@
Let $f(x)$ be a nonzero ${\mathcal{K}}$-CLC over a proper convex cone ${\mathcal{K}}$ and $v,x\in\operatorname{int}{\mathcal{K}}$. Then the sequence $\{f(x),D_{v}f(x),D_{v}^{2}f(x),\dots,D_{v}^{d}f(x)\}$ of positive real numbers forms a log-concave sequence.
@@
Since $f$ is ${\mathcal{K}}$-CLC, so it follows from Definition 3.12 that the univariate restriction $f(x+tv)$, for any $x,v\in\operatorname{int}{\mathcal{K}}$, is CLC as a polynomial in $t$ for all $t\geq 0$. According to Proposition 3.13, we know that all the coefficients of these univariate restricted polynomials are positive. Then the rest follows from Lemma 3.10. ∎
@@
The condition that the coefficients form a log-concave sequence is not sufficient, as the log-concavity of the univariate polynomial must hold for any $v\in{\mathbb{R}}^{n}$, not just for $v\in{\mathcal{K}}$, as required by Definition 3.12.

### v2

@@ 3.1. Proper Convex Cones
Using Lemma˜2.4 we show that $\mathcal{K}$}-nonnegativity is a sufficient condition for $\mathcal{K}$}-Lorentzian polynomials if $\mathcal{K}$} is contained in its dual cone, $\mathcal{K}$$^{\ast}$.
@@
Since $Q$ is $\mathcal{K}$}-nonnegative, so $Q({\mathcal{K}})\subseteq{\mathcal{K}}\subseteq{\mathcal{K}}^{\ast}$. Then the rest follows from Lemma˜2.4. ∎
Based on Lemma˜2.4 we show that $\mathcal{K}$}-nonnegativity is a necessary condition for $\mathcal{K}$}-Lorentzian polynomials if $\mathcal{K}$} contains its dual cone, $\mathcal{K}$$^{\ast}$.
@@
By Lemma˜2.4 we have the matrix $Q$ of $q$ has exactly one positive eigenvalue and $Q({\mathcal{K}})\subseteq{{\mathcal{K}}^{*}}$. Thus, $Q({\mathcal{K}})\subseteq{\mathcal{K}}^{*}\subseteq{\mathcal{K}}$. Therefore, $Q$ is $\mathcal{K}$}-nonnegative by Definition˜2.3. ∎
@@
The converse to Corollary˜3.2 is not true. It’s easy to construct counterexamples where ${\mathcal{K}}^{\ast}\subseteq{\mathcal{K}}$ and $Q({\mathcal{K}})\subseteq{\mathcal{K}}$ but $Q({\mathcal{K}})\not\subseteq{\mathcal{K}}^{\ast}$.
It’s shown that any closed convex cone contains its dual cone, i.e., ${\mathcal{K}}^{\ast}\subset{\mathcal{K}}$, if and only if there exist $y,z\in{\mathcal{K}}$ such that x=yzx=y-z and $\langle y,z\rangle=0$, cf. [Remarkcopositive] for detail. This is used to show that every $x\in{\mathbb{R}}^{n}$ has an (unique) orthogonal decomposition on $\mathcal{K}$} in [Cross] which have many well-known consequences. Additionally, the sum of two $\mathcal{K}$}-nonnegative matrices is $\mathcal{K}$} nonnegative. Furthermore, if $A$ and $B$ are $\mathcal{K}$}-nonnegative and $A$ is $\mathcal{K}$}-irreducible, then so is $A+B$, cf. [Bermannonnegative, Chap-1]. However, the sum of two log-concave polynomials may not be log-concave. The sufficient condition for the sum of two $\mathcal{K}$}-Lorentzian polynomials to also be $\mathcal{K}$}-Lorentzian can be found in [Cynthialog3, Lemma 3.3] for the nonnegative orthant, and [GPlorentzian, Theorem 4.15] for any proper convex cone $\mathcal{K}$}.
@@
[GPlorentzian, Theorem 4.15] Let $f,g\in{\mathbb{R}}^{d}_{n}[x]$ be two $\mathcal{K}$}-Lorentzian polynomials over the self-dual cone $\mathcal{K}$}. Then the sum $f+g$ is $\mathcal{K}$}-Lorentzian if there exist vectors $b,c\in{\mathcal{K}}$ such that $D_{b}f=D_{c}g\not\equiv 0$.
@@ 3.2. Perron-Frobenius Theorem for Cones
Let $\mathcal{K}$} be a proper convex cone. We can leverage the relationship between $\mathcal{K}$}-Lorentzian polynomials and the generalized Perron-Frobenius theorem [Kreinlinear], [Barkeralgebraic] to distinguish potential $\mathcal{K}$}-Lorentzian polynomial candidates from given log-concave polynomials. This also provides valuable insight into the structure of the cone K.
@@
(Generalization of Perron-Frobenius Theorem) [Kreinlinear], [Barkeralgebraic]: Let $A$ be a $\mathcal{K}$}-irreducible matrix with spectral radius $\rho$. Then
@@
By applying [GPlorentzian, Lemma 4.12] and [GPlorentzian, Prop:4.13] we derive the following characterization of the Hessian matrices evaluated at any point over the cone $\mathcal{K}$} for any $\mathcal{K}$}-Lorentzian polynomials where $\mathcal{K}$} is a proper convex cone.
@@
In special cases using the Definition˜2.2 and [GPlorentzian, Them:5.1] we have some necessary conditions for $\mathcal{K}$}-Lorentzian polynomials over the self-dual cone $\mathcal{K}$}.
@@
By Theorem˜3.5 we have more stronger necessary conditions for $\mathcal{K}$}-Lorentzian polynomials over the self-dual cone $\mathcal{K}$}.
@@
Since $f$ is a $\mathcal{K}$}-Lorentzian polynomial, so for any $a\in{\mathcal{K}}$, the nonsingular Hessian $H_{f}(a)$ is $\mathcal{K}$}-irreducible matrix and has exactly one positive eigenvalue. Therefore, if $H_{f}(a)$ is nonsingular, and the positive eigenvalue corresponds to the spectral radius, then by Theorem˜3.5, its corresponding eigenvector must lie in the interior of the cone $\mathcal{K}$}. ∎
The geometry of the set of symmetric Perron–Frobenius matrices has been studied in [SymmetricPFstructure]. In fact, it is not convex but is star convex with the identity matrix as the center. It’s a natural question to describe the geometry of the set of symmetric Perron–Frobenius matrices which have exactly one positive eigenvalue. We leave this question for future work.
@@ 3.3. Properties of $\mathcal{K}$}-CLC polynomials
Here we propose another definition for $\mathcal{K}$}-CLC multivariate polynomials by restricting them on certain lines. This allows us to show the log-concavity property among the coefficients of the corresponding univariate restrictions of a $\mathcal{K}$}-CLC polynomial. Recall that a univariate polynomial $f(t)$ is log-concave over an interval $I$ if $f(t)>0$ for all $t\in I$ and $f(t)f^{{}^{\prime\prime}}(t)\leq(f^{{}^{\prime}}(t))^{2}$ for all $t\in I$. Equivalently, a univariate polynomial $f(t)$ is log-concave over an interval $I$ if $f(t)>0$ for all $t\in I$ and $\frac{d^{2}}{dt^{2}}\log f(t)\leq 0$ for all $t\in I$. So, we can define $\mathcal{K}$}-CLC for any real multivariate polynomial as in Definition˜2.1, need not be homogeneous. If they are homogeneous, the set of $\mathcal{K}$}-Lorentzian polynomials of degree $d$ are equivalent to the set of $\mathcal{K}$}-CLC forms , cf. [GPlorentzian]). On the other hand, over the nonnegative orthant, a characterization for bivariate $\mathcal{K}$}-CLC form has been established in the literature, providing a resolution to Mason’s conjecture.
@@
[Gurvitsmultivariate], [Cynthialog3] $f=\sum_{k=0}^{n}c_{k}x^{n-k}y^{k}\in{\mathbb{R}}[x,y]$ is completely log- concave if and only if the sequence of nonnegative coefficients, $\{c_{0},\dots,c_{n}\}$ is ultra log-concave, i.e., for every $1<k<n$, $\left(\frac{c_{k}}{\binom{n}{k}}\right)^{2}\geq\frac{c_{k-1}}{\binom{n}{k-1}}\frac{c_{k+1}}{\binom{n}{k+1}}$.
@@
It’s straightforward to verify that $f(t)=\sum_{i=0}^{d}a_{i}t^{i}$ with $a_{i}>0$ is CLC for $t\geq 0$ if and only if its homogenization $f(t,u)=\frac{u^{d}}{d!}a_{0}+\frac{u^{d-1}}{(d-1)!}a_{1}t+\dots+\frac{u^{0}}{0!}a_{d}t^{d}$ is CLC over ${\mathbb{R}}^{2}_{\geq 0}$. Then the rest follows from Theorem˜3.9 and the identity $a_{i}^{2}(i!)^{2}\geq a_{i-1}a_{i+1}(i-1)!(i+1)!\Leftrightarrow ia_{i}^{2}\geq(i+1)a_{i-1}a_{i+1}$. ∎
@@
Due to Lemma˜3.10 and symmetry, $((d-k)!a_{k})^{2}\geq(d-k-1)!(d-k+1)!a_{k-1}a_{k+1}\Leftrightarrow(d-k)a_{k}^{2}\geq(d-k+1)a_{k-1}a_{k+1}\Leftrightarrow ia_{d-i}^{2}\geq(i+1)a_{d-i-1}a_{d-i+1}$. ∎
It is well known that convexity can be characterized in terms of one-dimensional convexity. More precisely, a function $f:{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}$ is convex over a convex set $\mathcal{K}$} if and only if for any $x\in{\mathcal{K}}$ and $v\in{\mathbb{R}}^{n}$, the function $g:{\mathbb{R}}\rightarrow{\mathbb{R}},g(t)=f(x+tv)$ is convex as function in $t$ for all $t$ such that $x+tv\in{\mathcal{K}}$ [Boydconvex]. In the same spirit, we introduce a line-restriction notion for log-concavity and complete log-concavity over cones.
Definition 3.12 (Linewise $\mathcal{K}$}-completely log-concave).
Let ${\mathcal{K}}\subseteq{\mathbb{R}}^{n}$ be a proper convex cone and let $f\in{\mathbb{R}}[x_{1},\dots,x_{n}]$. For $x\in\operatorname{int}{\mathcal{K}}$ and $v\in{\mathcal{K}}\setminus\{0\}$, set
$I_{x,v}\coloneqq\{\,t\in{\mathbb{R}}:x+tv\in\operatorname{int}{\mathcal{K}}\,\},\qquad g_{x,v}(t)\coloneqq f(x+tv),\quad t\in I_{x,v}.$
  • We say that $f$ is linewise log-concave over $\mathcal{K}$} if, for every $x\in\operatorname{int}{\mathcal{K}}$ and $v\in{\mathcal{K}}\setminus\{0\}$, the univariate function $g_{x,v}$ is log-concave on $I_{x,v}$.
  • We say that $f$ is linewise $\mathcal{K}$}-completely log-concave (linewise $\mathcal{K}$}-CLC) if, for every $x\in\operatorname{int}{\mathcal{K}}$ and $v\in{\mathcal{K}}\setminus\{0\}$, the univariate polynomial $g_{x,v}$ is completely log-concave on $I_{x,v}$ in the univariate sense (i.e., for every $m\geq 0$, the derivative $g_{x,v}^{(m)}$ is log-concave on $I_{x,v}$).
We next show that $\mathcal{K}$}-complete log-concavity in the sense of Definition˜2.1 implies the linewise property from Definition˜3.12.
Proposition 3.13.
Let ${\mathcal{K}}\subseteq{\mathbb{R}}^{n}$ be a proper convex cone, and suppose that $f\in{\mathbb{R}}[x_{1},\dots,x_{n}]$ is $\mathcal{K}$}-completely log-concave in the sense of Definition˜2.1. Then $f$ is linewise $\mathcal{K}$}-completely log-concave.
Proof.
Fix $x\in\operatorname{int}{\mathcal{K}}$ and $v\in{\mathcal{K}}\setminus\{0\}$, and let $I_{x,v}$ and $g_{x,v}$ be as in Definition˜3.12. We must show that $g_{x,v}$ is completely log-concave on $I_{x,v}$ in the univariate sense.
By Definition˜2.1, for every integer $m\geq 0$ the directional derivative
$D_{v}^{m}f\;=\;\underbrace{D_{v}\cdots D_{v}}_{m\ \text{times}}f$
is log-concave on $\operatorname{int}{\mathcal{K}}$, since each direction $v$ belongs to $\mathcal{K}$}. For $t\in I_{x,v}$ we have $x+tv\in\operatorname{int}{\mathcal{K}}$, and
$\frac{d^{m}}{dt^{m}}g_{x,v}(t)\;=\;\frac{d^{m}}{dt^{m}}f(x+tv)\;=\;D_{v}^{m}f(x+tv).$
Thus the $m$-th derivative of $g_{x,v}$ is the restriction of the log-concave function $D_{v}^{m}f$ (defined on $\operatorname{int}{\mathcal{K}}$) to the line segment $\{x+tv:t\in I_{x,v}\}$. Since restrictions of a log-concave function to a line are log-concave on their domain, it follows that $g_{x,v}^{(m)}$ is log-concave on $I_{x,v}$ for every $m\geq 0$.
This is precisely the definition of complete log-concavity for the univariate function $g_{x,v}$ on $I_{x,v}$, and therefore $f$ is linewise $\mathcal{K}$}-completely log-concave. ∎
@@
Let $f(x)$ be a strictly $\mathcal{K}$}-CLC over a proper convex cone $\mathcal{K}$}. Then for any $x,v\in{\mathcal{K}}$, the coefficients of $f(x+tv)$ are positive. In particular, if $f(x)$ is a strictly $\mathcal{K}$}-Lorentzian polynomial over a proper convex cone $\mathcal{K}$}, then for any $x,v\in{\mathcal{K}}$, the coefficients of $f(x+tv)$ are positive.
@@
It follows from Definition˜2.1 and Proposition˜3.14. ∎
@@
Let $K\subseteq{\mathbb{R}}^{n}$ be a proper convex cone and let $f(x)$ be a nonzero $K$-CLC polynomial. Let $d:=\deg f$ and let $x,v\in\operatorname{int}K$. Then the sequence
$\bigl\{f(x),\,D_{v}f(x),\,D_{v}^{2}f(x),\,\dots,\,D_{v}^{d}f(x)\bigr\}$
of positive real numbers is log-concave.
@@
Since $f$ is $K$-CLC in the sense of Definition˜2.1, it follows from Proposition˜3.13 that $f$ is linewise $K$-CLC in the sense of Definition˜3.12. Hence, for any $x,v\in\operatorname{int}K$, the univariate restriction
$g(t)\coloneqq f(x+tv)$
is completely log-concave on $t\geq 0$.
By Proposition 3.14, for any $x,v\in\operatorname{int}K$ all coefficients of $g(t)$ are positive. Writing
$g(t)=f(x+tv)=f(x)+tD_{v}f(x)+\frac{t^{2}}{2!}D_{v}^{2}f(x)+\cdots+\frac{t^{d}}{d!}D_{v}^{d}f(x),$
we see that the coefficients are
$a_{k}=\frac{1}{k!}D_{v}^{k}f(x),\qquad k=0,1,\dots,d.$
Since $g$ is univariate CLC with positive coefficients, Lemma 3.10 applies and implies that the sequence
$\{0!a_{0},\,1!a_{1},\,\dots,\,d!a_{d}\}$
is log-concave. Substituting $a_{k}=\frac{1}{k!}D_{v}^{k}f(x)$, we obtain
$0!a_{0}=f(x),\quad 1!a_{1}=D_{v}f(x),\quad\dots,\quad d!a_{d}=D_{v}^{d}f(x),$
so the sequence
$\bigl\{f(x),\,D_{v}f(x),\,D_{v}^{2}f(x),\,\dots,\,D_{v}^{d}f(x)\bigr\}$
is log-concave, as claimed. ∎
@@
The converse of Theorem˜3.16 does not hold. Even if, for every $x\in\operatorname{int}K$ and every $v\in\operatorname{int}K$, the sequence
$\bigl\{f(x),D_{v}f(x),\dots,D_{v}^{d}f(x)\bigr\}$
is log-concave, this does not imply that $f$ is $K$-CLC.